OpenStudy (anonymous):

help! simplify {(y^2 + y)/(3xy)}*{(3x + 3)/(y^2+2y+1)}

OpenStudy (anonymous):

\[\left( \frac{ y^2+y }{ 3xy } \right)\left( \frac{ 3x+3 }{ y^2+2y+1 } \right)\] \[\left(\frac{ y(y+1) }{ 3xy } \right)\left( \frac{ 3(x+1) }{ (y+1)^2 } \right)\] \[\left( \frac{ x+1 }{ x(y+1) } \right)\]

OpenStudy (anonymous):

thank you!!!